OpenStudy (loser66):

Need help understanding problem: Let S be a subset of a set X. Let R be the ring of real-valued functions on X, and let I be the set of real-valued functions on X whose restriction to S is zero. Show that I is an ideal in R. Please, help

OpenStudy (loser66):

I don't understand : " WHOSE RESTRICTION TO S IS ZERO", what does it mean? This is my interpret: \(S\subset X\) , \(Fun(X, R)\) , \(I\subset Fun(X,R)\)

OpenStudy (loser66):

If \(f\in I\), then \(\forall x \in S, f(x) =0\) . Am I right?

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

the second statement looks right to me, in other words L is the set of functions that vanish on S

OpenStudy (misty1212):

should be pretty clearly an ideal, since if \(f\in I , g\in R\) then \(fg\in I \)

OpenStudy (loser66):

You mean S is kernel of Fun?

OpenStudy (loser66):

nope, f(S) is kernel, right?

OpenStudy (misty1212):

i think you lost me there R is the ring right?

OpenStudy (loser66):

yes

OpenStudy (misty1212):

it is the ring of real valued function on X I is a subset of R, it is the set of functions that vanish on S

OpenStudy (loser66):

I confused, can you visualize it? Like 

OpenStudy (misty1212):

that is not the ring, that is just the set on which the functions act

OpenStudy (misty1212):

say X is the interval [-1,1] and S = {0} then I would be the set of functions \(f\) on [-1,1] for which \(f(0)=0\)

OpenStudy (loser66):

I think I got it :) Let me interpret it and you check, ok?

OpenStudy (misty1212):

the ring is the set of all real valued functions on X, I is the set of all real valued function on X that are zero on S what you have to show is the following a) (I,+) is a subgroup of R, i.e. if you add to functions in I you get another function in I etc b) \(\forall f\in I, g\in R, fg\in I\)

OpenStudy (loser66):

R = {set of real-valued functions on X} =\(\{f_a, f_b, f_c, ........\}\) and \(I\subset R\) such that \(f_I (S) =0\)

OpenStudy (misty1212):

i wouldn't bother to list them with indices no one says they are countable, and you don't reallyl want to introduce some uncountable index set. just say R is what it is, the set of all real valued functions on X

OpenStudy (misty1212):

in fact they are uncountable if X is, leave that part alone

OpenStudy (loser66):

So, to prove I is an ideal, I have to prove it is a subring first, right? then the properties of left/right ideal to get it is an ideal, right?

OpenStudy (misty1212):

\[I=\{f\in R:\forall s\in S, f(s)=0\}\]

OpenStudy (loser66):

The reason I list R out is to easily define identities

OpenStudy (misty1212):

oh no don't list anything just pick two arbitrary ones \[f,g\in I\implies f+g\in I \] etc

OpenStudy (loser66):

Thanks a lot. I will show my stuff, please check then. :)

OpenStudy (misty1212):

ok the point of this exercise is to see if you know exactly what you have to show to show a subset of a ring is an ideal in the ring. list all the steps showing them in this case should not be hard at all, the hard part is knowing what to show

OpenStudy (loser66):

1) I is closed under addition and additive inverse: Let f, g in I, then f(s) =0 and g(s) =0 f(s) + g(s) = (f+g)(s) = 0+0=0 --> (f+g) in I Moreover, f(s) + g(s) =0 --> f(s)= -g(s) That is g(s) is additive inverse of f(s) Hence, 1) follows 2) I is closed under multiplication f(s)*g(s) = (f*g)(s) =0--> f*g is in I Hence 2) follows 1), 2) --> I is a subring of R

OpenStudy (loser66):

Now, prove I is a left ideal let g in R , \(g(s) \neq 0\), f is I then g(s)f(s) = (gf)(s) =0 , hence g(s)f(s) \(\in I\) similarly for right ideal we get I is ideal Am I right?

OpenStudy (misty1212):

looks reasonable to me i always forget whether you have to show that 0 is in it, but of course it is if you want to be a bit more precise at the beginning you could say " let \(f,g\in I\) then \(\forall s\in S, f(s)-g(s)=0\implies f-g\in I\)"

OpenStudy (misty1212):

as for the "left, right" business, multiplication is commutative, so the is a commutative right, not necessary

OpenStudy (loser66):

In my book, they say a ring is abelian group, hence it automatically has additive identity 0

OpenStudy (misty1212):

yes what i meant was you may need to show that 0 is in the set, but maybe not if \(f-g\) is in it then \(f-f=0\) is in it, so that step is probably unnecessary

OpenStudy (loser66):

:) Again, thanks a ton.